A study analyzed data from the National Longitudinal Study of Adolescent Type numbers in the boxes. Health. Participants were followed into adulthood. Fach study participant was 7 Paits ponts categorized as to whether they were obese (BMI -30) or not and whether they art a z pothts were dating, cohabiting, or married. The researchers were trying to determine the effect of relationship status on obesity The table below 6 points summarizes thePnts results: Dating Cohabiting Mamed Total Obese 81 104 2 Not obese 35932S Total 440429 Caleulate the column conditional percentages for participants who were Married: (Important: Input percentages and round to the nearest : decimal place) Married Obese 134.7 % Not obese Perc 2 Total Total Pere Total Perc
Answer #1

Perc 2 = P(Not obese| married) = P(No obese and married)/P(married)

= 277/424

= 0.6533

= 65.3 %

total perc = 100